OpenStudy (anonymous):

What is the equation of the line in slope-intercept form that passes through the point (1, 3) and is perpendicular to the line y = –x + 2? y = x – 4 y = x + 2 y = –x + 2 y = –x – 4

OpenStudy (anonymous):

@Valkarie70

OpenStudy (anonymous):

slope intercept is the same as Y=Mx+b, so now we must find the equations that are +x instead of -x in them for them to be perpendicular, (this determines the direction of the slope). C and D are out of the picture because they are negativ or parrallel and you want perpendicular, so graph A and B to find which one passes through (1,3) Answer: B) y=x+2

OpenStudy (anonymous):

You use the negative reciprocal for the new slope, so, the original slope is -1 and the new slope is 1. Equation is in the form y = mx + b. We know slope, so y = x + b. Substitute in the point to get "b": 3 = 1 + b, so b = 2 and m = 1.